# DualReal: Adaptive Joint Training for Lossless Identity-Motion Fusion in Video Customization

### v1

@@ 4 Experiment
CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ T.Fli $\uparrow$ T.Cons $\uparrow$ MS $\uparrow$ Dynamic Degree
@@
Table 1: Quantitative comparison of personalization video generation for customized subject and motion. We highlight the best and second-best values for each metric. “T.Cons” and “T.Flickering” denotes Temporal Consistency and Temporal Flickering, respectively. Compared with other methods, DualReal achieved average improvements of 21.7% on CLIP-I and 31.8% on DINO-I, recorded the best results on three motion quality metrics (T.Cons, Motion Smoothness(MS), and Temporal Flickering), and ranked second on CLIP-T. The motion datasets achieve an average Dynamic Degree of 12.02 and parenthetical values quantify the current method’s deviation from this benchmark to determine the intensity consistency of movement.
@@ 4.1 Setup
Datasets. The evaluation datasets are divided into two components: identity images and motion videos. For identity customization, 50 subjects are strictly selected from previous works  [23, 24] and Internet collections (including pet, plush, etc.), with each subject containing 3–10 images. For motion customization, 21 motion sequences with challenging dynamic patterns are collected from public datasets [41, 40]. Additionally, each case is provided with 50 various prompts containing different editability (i.e., decoration or environment) to sequentially evaluate the method’s editability and scene versatility.
Baselines. Among existing methods, only DreamVideo [48] achieves customization of both subject and motion. For fair comparison, we implement two approaches from the same backbone: (1) CogVideoX-5B [52]: Sequential full-parameter fine-tuning with identity then motion data as in DreamBooth [34] paradigm. (2) LoRA fine-tuning [17]: Separate training of two LoRA modules for identity and motion, then fuse their parameters during inference. Additionally, the identity module of MotionBooth [49] introduces irrelevant random videos during training to preserve the model’s motion capability, so we compare our approach with this method as well. In summary, we evaluate our results against DreamVideo, CogVideoX-5B, LoRA fine-tuning and MotionBooth to provide a more comprehensive performance analysis.
@@ 4.2 Main Results
Quantitative results. As shown in Tab. 1, DualReal achieved average improvements of 21.7% on CLIP-I and 31.8% on DINO-I, recorded the best results on three motion quality metrics (T.Cons, Motion Smoothness, and Temporal Flickering), and ranked second on CLIP-T. Although our DD metric for quantifying motion intensity is not high, we evaluate all motion data and found an average DD of 12.02. Our metric deviates slightly from it, proving there is no collapse in motion amplitude. Overall, our method significantly enhances motion coherence and identity fidelity while preserving text consistency, further validating our adaptive joint training approach.
@@ 4.3 Ablation Studies
Table 2: Quantitative ablation studies on each component. We implement Dual-aware Adaptation removal by separately training the two modalities and directly blends their parameters during inference, following the approach of DreamVideo. The motion datasets achieve an average Dynamic Degree of 9.04.
Figure 6: Qualitative ablations studies on each component. (1) Omitting Dual-aware Adaptation introduces artifacts on the subject’s hands and chin, significantly reducing clarity. Using fixed weights for the dimensional adapters without the StageBlender Controller causes the hands to become overly adapted to the motion pattern, and removing weight grouping reduces identity fidelity and background detail.
@@
Quantitative experiment on each component. The ablation study in Tab. 2 shows that removing Dual-aware Adaptation or the StageBlender Controller slightly increases the text consistency metric CLIP-T but significantly decreases identity similarity and motion intensity, highlighting the need for joint dimension training and granular control. Additionally, metric changes diminish slightly when weight groups are removed.
Qualitative experiment on each component. Qualitative results in Fig. 6 show that omitting Dual-aware Adaptation introduces artifacts on the subject’s hands and chin, significantly reducing clarity. Using fixed weights for the dimensional adapters without the StageBlender Controller (i.e., direct fusion at inference) causes the hands to become overly adapted to the motion pattern. Removing weight grouping (i.e., uniformly modulating all blocks) reduces identity fidelity and background detail compared to our final configuration. These findings confirm that each component is essential for high-quality customization generation.
@@ 4.4 Visual analysis of StageBlender Controller
As Fig. 7 shows, shallow blocks (Groups 1–6, blue) progressively increase identity weights during denoising, emphasizing early identity preservation. In contrast, the deepest block (Group 7, red) steadily raises motion weights to enhance motion modeling. Overall (orange dashed line), as denoising advances, the model increasingly prioritizes identity preservation over motion generation, highlighting the distinct roles of different network depths.
These observations further confirm that: (1) as denoising progresses, the diffusion model’s emphasis shifts monotonically between identity and motion, with a growing focus on identity; and (2) DiT networks of different depths divide the tasks of modeling identity and motion differently at each denoising step, with the deepest network focusing on motion patterns and showing increased enhancement as denoising advances.
@@ 5 Conclusion
In this paper, we propose DualReal, a novel approach for customized video generation given a subject and motion. DualReal adaptively trains identity and motion jointly, resolving dimensional conflicts and enabling universal sample customization. Our framework leverages the prior from one dimension to guide the training of the other, while preventing information leakage through a regularization strategy. Simultaneously, we use a controller to guide the lossless fusion of modes based on various denoising stages and DiT depths. Evaluated on a more comprehensive evaluation benchmark, our method improves CLIP-I and DINO-I metrics by 21.7% and 31.8% on average, and achieves top performance on nearly all motion quality metrics, demonstrating the efficiency of our adaptive joint training framework.

### v2

@@ 4 Experiment
CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ T.Flickering $\uparrow$ T.Cons $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree
@@
Table 1: Quantitative comparison of personalization video generation for customized subject and motion. We highlight the best and second-best values for each metric. “T.Cons" and “T.Flickering" denote Temporal Consistency and Temporal Flickering, respectively. Compared with other methods, DualReal achieved average improvements of 21.7% on CLIP-I and 31.8% on DINO-I, recorded the best results on three motion quality metrics (T.Cons, Motion Smoothness, and Temporal Flickering), and ranked second on CLIP-T. The motion datasets achieve an average Dynamic Degree of 12.02, and parenthetical values quantify the current method’s deviation from this benchmark to determine the intensity consistency of movement.
@@ 4.1 Setup
Datasets. The evaluation datasets are divided into two components: identity images and motion videos. For identity customization, 50 subjects are strictly selected from previous works [26, 27] and Internet collections (including pet, plush, etc.), with each subject containing 3–10 images. For motion customization, 21 motion sequences with challenging dynamic patterns are collected from public datasets [45, 44]. Additionally, each case is provided with 50 various prompts containing different editability (i.e., decoration or environment) to evaluate the method’s editability and scene versatility sequentially.
Baselines. Among existing methods, DreamVideo [52] achieves customization of both identity and motion. For fair comparison, we implement two approaches from the same DiT backbone: (1) CogVideoX-5B [56]: Sequential full-parameter fine-tuning with identity then motion data as in the DreamBooth [38] paradigm. (2) LoRA fine-tuning [18]: Separate training of two LoRA modules for identity and motion, then fuse their parameters during inference. Additionally, the identity module of MotionBooth [53] introduces irrelevant random videos during training to preserve the model’s motion capability, so we compare our approach with this method as well. In summary, we evaluate our results against DreamVideo, CogVideoX-5B, LoRA fine-tuning, and MotionBooth to provide a more comprehensive performance analysis.
@@ 4.2 Main Results
Quantitative results. As shown in Tab. 1, DualReal achieved average improvements of 21.7% on CLIP-I and 31.8% on DINO-I, recorded the best results on three motion quality metrics (T.Cons, Motion Smoothness, and Temporal Flickering), and ranked second on CLIP-T. Although our DD metric for quantifying motion intensity is not high, we evaluated all motion data and found an average DD of 12.02. Our metric deviates slightly from it, proving there is no collapse in motion amplitude. Overall, our method significantly enhances motion coherence and identity fidelity while preserving text consistency, further validating our adaptive joint training approach.
@@ 4.3 Ablation Studies
Table 2: Quantitative ablation studies on each component. We implement Dual-aware Adaptation removal by separately training the two modalities and directly blending their parameters during inference, following the approach of DreamVideo. The motion datasets achieve an average Dynamic Degree of 9.04.
Figure 6: Qualitative ablation studies on each component. (1) Omitting Dual-aware Adaptation introduces artifacts on the subject’s hands and chin, significantly reducing clarity. Using fixed weights for the dimensional adapters without the StageBlender Controller causes the hands to become overly adapted to the motion pattern, and removing weight grouping reduces identity fidelity and background detail.
@@
Quantitative experiment. The ablation study in Tab. 2 shows that removing Dual-aware Adaptation or the StageBlender Controller slightly increases the text consistency metric CLIP-T but significantly decreases identity similarity and motion intensity, highlighting the need for joint dimension training and granular control. Additionally, metric changes diminish slightly when weight groups are removed.
Qualitative experiment. Qualitative results in Fig. 6 reveal that omitting Dual-aware Adaptation produces artifacts on the hands and chin, degrading clarity. Using fixed weights for the dimensional adapters without the StageBlender Controller (i.e., direct fusion at inference) overfits the hands to motion patterns. Removing weight grouping (i.e., uniform block modulation) weakens identity fidelity and background detail. These observations confirm that every component is essential for high-quality customized generation.
@@ 4.4 Visual analysis of StageBlender Controller
As Fig. 7 shows, shallow blocks (Groups 1–6, blue) progressively increase identity weights during denoising, emphasizing early identity preservation. In contrast, the deepest block (Group 7, red) steadily raises motion weights to enhance motion modeling. Overall (orange dashed line), as denoising advances, the model increasingly prioritizes identity preservation over motion generation, highlighting the distinct roles of different network depths. These observations further confirm that: (1) as denoising progresses, the diffusion model’s emphasis shifts monotonically between identity and motion, with a growing focus on identity; (2) DiT networks of different depths divide the tasks of modeling identity and motion differently at each denoising step, with the deepest network focusing on motion patterns and showing increased enhancement as denoising advances.
@@ 5 Conclusion
In this paper, we propose DualReal, a novel approach for customized video generation given a subject and motion. DualReal adaptively trains identity and motion jointly, resolving dimensional conflicts and enabling universal sample customization. Our framework leverages the prior from one dimension to guide the training of the other, while preventing information leakage through a regularization strategy. Simultaneously, we use a controller to guide the high-fidelity fusion of modes based on various denoising stages and DiT depths. Evaluated on a more comprehensive evaluation benchmark, our method improves CLIP-I and DINO-I metrics by 21.7% and 31.8%, and achieves top performance on nearly all motion quality metrics, demonstrating the efficiency of our adaptive joint training framework.
6 Acknowledgment
This research is supported by Artificial Intelligence National Science and Technology Major Project 2023ZD0121200, and National Natural Science Foundation of China under Grant 62222212 and 623B2094.